OpenStudy (anonymous):

Carol and Nora left the library at the same time. Carol biked east at a speed of 8 miles per hour. Nora biked west at a speed of 10 miles per hour. How far apart were Carol and Nora after 3 hours?

OpenStudy (aaronandyson):

D=RT D=8*3+10*3 D=24+30 D=54 MILES

OpenStudy (anonymous):

thanks here is a medal and i will fan you